lazyking4life:

Given m \| nm∥n, find the value of x. (2x+9)=(7x+24)

Hoodmemes:

I'll try and answer this (Warning: I'm not good at math like this is so don't 100% trust my answer, and I'm getting help from a sibling) I (think) it goes like this : We can put together the variable terms on one side of the (equation) and then put together constraint terms on another side. \[(2x-9)+(-7x+9)=(7x-24)+(-7x+9)\] I'm thinking we can take away the "parentheses"- "()" now So it would be like this: \[2x-9+-7x+9=7x-24+-7x+9\] We now put the expressions in groups of like terms to make em easy I'll try and answer this (Warning: I'm not good at math like this is so don't 100% trust my answer, and I'm getting help from a sibling) I (think) it goes like this : We can put together the variable terms on one side of the (equation) and then put together constant terms on another side. \[(2x-9)+(-7x+9)=(7x-24)+(-7x+9)\] I'm thinking we can take away the "parentheses"- "()" now So it would be like this: \[2x-9+-7x+9=7x-24+-7x+9\] We now put the expressions in groups of like terms to make em easy to solve so it would be like : \[2x-7x-9+9=7x-7x-24+9\] I'd say solve this side first (2x-7x-9+9) I think how u do it is that u subtract 2 and 7 \[2x-7x=-5x\] and now we solve (7x-7x-24+9) \[7x-7x-24+9\rightarrow-24+9\] \[-24+9=-15\] We can divide both (-5x) and (-15) by (5) So \[-15\div5=-3\] take (-3) and make it (3) and take (-5x) and turn it into (5x) \[5x \div5= x\] So I think x=3 (Now this could be wrong so don't fully trust my answer, hope this helps,if it is wrong can somebody explain to me how to do this? )